OpenStudy (anonymous):

If EF= 2x-18, FG=3x-13, and EG=29 find the values of x, EF, and FG

OpenStudy (anonymous):

@McDonaldsGirl<3, @TheRealMeeeee

OpenStudy (anonymous):

I don't know this kind of math sorry

OpenStudy (therealmeeeee):

What's the choices

OpenStudy (anonymous):

oh ok thanks anyway @McDonaldsGirl<3

OpenStudy (anonymous):

Answer chocies: a.x=5, EF=-8,FG=2 b. x=12, EF=6, FG=23 c. x=12, EF=42,FG=49 d. x=5,EF=6, FG=23

OpenStudy (anonymous):

@Abbles

OpenStudy (therealmeeeee):

EF + FG = EG (2x - 12) + (3x - 15) = 23 5x - 27 = 23 5x = 50 x = 10 EF = 2x - 12 EF = 2(10) - 12 EF = 20 - 12 EF = 8 FG = 3x - 15 FG = 3(10) - 15 FG = 30 - 15 FG = 15

OpenStudy (anonymous):

Thanks @TheRealMeeeee

OpenStudy (therealmeeeee):

np

OpenStudy (abbles):

So 2x-18 + 3x-13 = 29 5x - 31 = 29 5x = 60 x = 12 Then plug in 12. EF= 2x-18 EF = 2(12) - 18 EF = 24 - 18 EF = 6 FG=3x-13 FG = 3(12) - 13 FG = 36-13 FG = 23

OpenStudy (abbles):

@TheRealMeeeee I think you used wrong information... EF is 2x - 18 not 2x - 12

OpenStudy (therealmeeeee):

Oh I do

OpenStudy (anonymous):

thanka @Abbles you are a big help

OpenStudy (anonymous):

thanks*

OpenStudy (abbles):

anytime :)